how many years will it take an investment to triple itself if the interest rate is 9% compounded annually?
Answer by mananth(16949)

Principal P = 100
Amount= 300
years=n
compounded (t) 1
Rate = 9
n = ln(FV / PV) / ln(1 + r)

n=(ln( 300/100 )/ln(1+0.09)
n=(1.1)/ln1.09
n=( 1.1 / 0.09
n= 12.75 years